# Neural Approaches for Multi-Objective Routing on Multigraphs

### v2

@@ Abstract
Learning-based methods for routing have gained significant attention in recent years, both in single-objective and multi-objective contexts. Yet, existing methods are unsuitable for routing on multigraphs, which feature multiple edges with distinct attributes between node pairs, despite their strong relevance in real-world scenarios. In this paper, we propose two graph neural network-based methods to address multi-objective routing on multigraphs. Our first approach operates directly on the multigraph by autoregressively selecting edges until a tour is completed. The second model first simplifies the multigraph via a learned pruning strategy and then performs routing on the resulting simple graph. We evaluate both models empirically and demonstrate their strong performance across a range of problems and distributions [^1] Our code will be released publicly upon paper acceptance..
@@ 1 Introduction
Routing problems, including the well-known Traveling Salesman Problem (TSP), are central for optimizing operations in modern transport systems. A key assumption is often that an optimal path can be selected beforehand between each pair of locations. However, in reality, several paths may exist that cannot be chosen between a priori. This naturally leads to a multigraph representation. Notably, by planning on a multigraph rather than a simple graph, Ben Ticha et al. (2017) demonstrated a cost reduction of up to $10.5\%$ for the vehicle routing problem with time-windows. Similarly Lai, Caliskan Demirag, and Leung (2016) demonstrated average savings of circa $5\%$ for a heterogeneous vehicle routing problem. Such results can be pivotal in for instance last-mile-delivery, which accounts for $41\%$ of total supply chain cost according to Capgemini Research Institute (2023).
Yet, routing problems on simple graphs are already difficult to solve, especially in real-time, as they are generally NP-hard. The multigraph setup further increases this complexity. Moreover, such a framework is particularly interesting in the Multi-Objective (MO) case, where each edge may be associated with several conflicting cost criteria, such as time, distance or energy. MO routing is typically harder, as the number of Pareto optimal solutions can grow exponentially with the size of the problem (Ehrgott and Gandibleux 2000). Clearly, MO routing on multigraphs poses significant challenges, while remaining highly relevant for real-world applications.
Recently, however, learning-based methods have emerged as promising options for routing, offering speed, accuracy, and broad applicability across problems (Kool, van Hoof, and Welling 2018; Drakulic, Michel, and Andreoli 2025). Yet, current approaches are designed for simple graphs and not suited for the multigraph representation. This is especially true in the MO setting, where neural approaches have demonstrated excellent results on Euclidean instances (Lin, Yang, and Zhang 2022; Fan et al. 2025), but lack support even for asymmetric graphs, in spite of their higher practical relevance (Boyacı, Dang, and Letchford 2021).
We attempt to bridge the gap between learning-based methods for MO routing and accurate network representations. Concretely, our main contributions are:
  • We design models to handle multigraph inputs, where not only the next node but also the edge must be selected. To our knowledge, we present the first neural solvers designed for such input structures. Although we deal with MO optimization, our approach can be highly useful in single-objective cases too.
  • Our models are also capable of handling a variety of MO routing problems on asymmetric graphs. In contrast, current neural MO methods are either problem-specific or designed exclusively for the Euclidean setting, where edge costs are symmetric and determined only by the node coordinates.
@@
  • Experimentally, we show that our approach achieves competitive results for several MO routing problems on asymmetric graphs and multigraphs.
@@ 2.1 Learning for Routing
Motivated by the success of end-to-end learning in fields such as image classification and language modeling, neural solutions for routing have gained increasing attention during the last years. Broadly, methods can be classified by how they obtain tours. Two prominent alternatives are autoregressive construction, based on iterative building of tours one node at a time, and non-autoregressive construction, that outputs heatmaps in a one-shot fashion for down-stream decoding with some search method. Examples from the first category include Vinyals, Fortunato, and Jaitly (2015); Kool, van Hoof, and Welling (2018); Kwon et al. (2020). Examples from the second category are Joshi, Laurent, and Bresson (2019); Fu, Qiu, and Zha (2021); Sun and Yang (2023). For a more comprehensive discussion of these methods, we refer the reader to Appendix A, which presents additional related work.
The literature on MO routing is more sparse than in the single-objective case. A key contribution was made by Lin, Yang, and Zhang (2022), who proposed an autoregressive construction approach using a single neural model to learn the Pareto set of solutions. Regarding asymmetric problems, the only existing learning-based methods we are aware of are those introduced by Santiyuda et al. (2024) and Zhou, Kong, and Wang (2025). However, both these approaches are tailored to specific problems and have been evaluated within narrowly defined domains. In contrast, our method is capable of addressing a broad class of routing problems and is evaluated across various problems and settings.
@@ 2.2 Routing on Multigraphs
The multigraph representation has become more and more common for MO shortest path problems during recent years (Weiszer, Burke, and Chen 2020; Beke, Weiszer, and Chen 2021). Such a framework has also been applied to combinatorial routing problems like Vehicle Routing Problems (VRPs), but to a lower extent. Garaix et al. (2010) performed the first study in this regard and later works include Lai, Caliskan Demirag, and Leung (2016); Ben Ticha et al. (2017); Tikani, Setak, and Demir (2021).
A common theme across these studies is that the multigraph representation leads to considerably improved solutions, while simultaneously increasing the complexity of these already difficult problems. Consequently, learning-based methods should be able to fill a clear need for new solvers in this area, as they have proven to be efficient, effective and broadly applicable. To the best of our knowledge, this work presents the first learning-based approach to address combinatorial routing problems such as the TSP and VRPs on multigraphs.
@@ 3 Problem Formulation
where $f:\Pi\to\mathbb{R}^{m}$ with $m>1$. Here, $\pi\in\Pi$ represents a feasible route in the multigraph $G$, encoded as a sequence of edges, and $f(\pi)$ is the associated cost vector. Due to the combinatorial nature of the feasible solution space $\Pi$ and the multiple conflicting objectives, the problem is generally NP-hard.
@@
To obtain this set, we utilize a typical method: decomposition through scalarization. The idea is to translate the MO problem into several subproblems that can be solved separately. Let $\lambda\in\mathbb{R}^{m}$ such that $\lambda_{i}\geq 0\>\forall i$ and $\sum_{i}\lambda_{i}=1$ be a vector that controls the preference among the objectives. Then, under Linear Scalarization (LS), the single-objective cost is $f_{\lambda}(\pi)=\sum_{i}\lambda_{i}f_{i}(\pi)$ and the corresponding subproblem is $\min_{\pi}f_{\lambda}(\pi)$. Unfortunately, it is well-known that there may be solutions in the PS that do not solve an LS subproblem for any preference $\lambda$ (Ehrgott 2005). The Chebyshev scalarization has more favorable properties. Let $z^{*}_{i}$ be an ideal value for objective $i$. Then, the Chebyshev scalarized cost is
@@
Given a solution in the PS, there always exists a preference $\lambda$ and corresponding Chebyshev subproblem for which the solution is optimal (Choo and Atkins 1983).
@@ 4 GNN-Based Multigraph Solver
Here, ||||| | is the concatenation operation, $e_{l}$ denotes an edge embedding for edge $l$, $W_{1}^{\prime}$ and $W_{1}^{\prime\prime}$ are trainable weight matrices and $E^{+}(i)$, $E^{-}(i)$ denote outgoing and incoming edges respectively. New edge embeddings are then computed according to
@@
where $\text{start}(l)$ and $\text{end}(l)$ denote the start and end nodes of the directed edge $l$. The attention layers are arranged into a single GREAT-layer together with feed-forward layers, residual connections and normalization according to the original transformer architecture (Vaswani et al. 2017). Note that the residual connections are important as they allow parallel edges to have differing embeddings.
@@ 4.1 Edge-Based GMS
$p_{\theta(\lambda)}(\pi\mid s)=\prod_{t=1}^{T}\;p_{\theta(\lambda)}(\pi_{t}\mid\pi_{1:t-1},s).$ (9)
Here, $\theta(\lambda)$ is the weights of the model given the current preference $\lambda$. As Navon et al. (2021) and Lin, Yang, and Zhang (2022), we utilize a Multi-Layer Perceptron (MLP) hyper-network to preference-condition the decoder, while keeping the encoder preference-agnostic. Consequently we form the decoder weights according to $\theta_{\text{dec}}(\lambda)=\text{MLP}(\lambda)$ and set
$\theta(\lambda)=[\theta_{\text{enc}},\theta_{\text{dec}}(\lambda)].$ (10)
Regarding the decoder architecture, it is an edge-based variant of the Multi-Pointer (MP) decoder from Jin et al. (2023). We detail this component further in Appendix B.
Note that a multigraph with $N$ nodes and $M$ edges between each pair of nodes has $\mathcal{O}(MN^{2})$ edges in total. Moreover, the decoder must be run $\mathcal{O}(N)$ times to construct a full tour and we roll out $\mathcal{O}(N)$ sample trajectories according to the POMO framework (Kwon et al. 2020). Hence, an edge-based decoder scales as $\mathcal{O}(MN^{4})$ compared to a node-based one that scales as $\mathcal{O}(N^{3})$. Consequently, we propose a second architecture that employs a node-based decoder and thereby scales better to higher node counts in terms of memory requirements and runtime.
@@ 4.2 Dual Head GMS
Again, the weights $\tilde{\theta}(\lambda)$ are generated by an MLP hyper-network. We detail the architecture for the selection-decoder further in Appendix B.
The edge choice from the first decoder is fed back into the encoder, where the embeddings of corresponding edges are aggregated into node embeddings using (7). After this final GREAT layer, we find it beneficial to add a small number of standard transformer layers to obtain more expressive node embeddings for the second decoder, which constructs routes.
The routing-decoder is a preference-conditioned version of the multi-pointer decoder (Jin et al. 2023). Similarly to the edge-based version above, given the partial tour $\pi_{1:t-1}$, this decoder outputs the probabilities $p_{\theta(\lambda)}(\pi_{t}\mid\pi_{1:t-1},s,\mathcal{E})$ for the next construction step, where $\pi_{t}$ now denotes a node and the probability depends on the edge choice $\mathcal{E}$.
@@
Figure 2: Dual Head GMS
@@ 4.3 Training Algorithm
Our first proposed model, GMS-EB, is trained using the Multi-Objective REINFORCE from Lin, Yang, and Zhang (2022). Given a preference $\lambda\sim\Lambda$ and instance $s\sim S$, the model parameterizes a policy $p_{\theta(\lambda)}(\cdot\mid s)$ according to (9). A tour $\pi$ is associated with a Chebyshev reward corresponding to the negative scalarized cost (5):
@@
To stabilize the training, we form a baseline $b_{\lambda}(s_{i})$ by rolling out $K$ tours for the same instance $s_{i}$ and averaging the reward, according to the POMO framework (Kwon et al. 2020). The policy gradient is then utilized to update the model weights and it is approximated, given one preference $\lambda$, according to
$\nabla J(\theta)\approx\frac{1}{BK}\sum_{i,j=1}^{B,K}A_{ij}\nabla_{\theta}\log(p_{\theta(\lambda)}(\pi_{ij}\mid s_{i})),$ (14)
where $B$ is the batch size of instances and the advantage is
$A_{ij}=R_{\lambda}(\pi_{ij})-b_{\lambda}(s_{i}).$ (15)
For the second model, the training is more involved. Given an edge selection $\mathcal{E}\sim q_{\tilde{\theta}(\lambda)}$ from the selection-head according to (11), the routing-head parameterizes a policy $p_{\theta(\lambda)}(\cdot\mid s,\mathcal{E})$ and obtains a scalar reward $R^{(2)}_{\lambda}(\pi)=R_{\lambda}(\pi)$ by (12) as usual. With a fixed $\tilde{\theta}$, the objective for this head is to maximize
@@
On the other hand, the task of the selection-head is to ensure that the edge choice is optimal. Let $\tilde{\pi}(\mathcal{E})$ denote the optimal tour on the pruned graph given selection $\mathcal{E}$. Then, define the reward and objective function based on the optimal tour according to
$\begin{split}&R^{(1)}_{\lambda}(\mathcal{E})=R_{\lambda}(\tilde{\pi}(\mathcal{E})),\\ &J_{1}(\tilde{\theta})=\textup{E}_{\lambda\sim\Lambda,\,s\sim S,\,\mathcal{E}\sim q_{\tilde{\theta}}}[R^{(1)}_{\lambda}(\mathcal{E})].\end{split}$ (17)
@@
$\begin{split}&\nabla J_{1}(\tilde{\theta})\approx\frac{1}{BK_{1}}\sum_{i,j=1}^{B,K_{1}}A^{(1)}_{ij}\nabla_{\tilde{\theta}}\log(q_{\tilde{\theta}}(\mathcal{E}_{ij}\mid s_{i})),\\ &\nabla J_{2}(\theta)\approx\frac{1}{BK_{1}K_{2}}\sum_{i,j,k=1}^{B,K_{1},K_{2}}A^{(2)}_{ijk}\nabla_{\theta}\log(p_{\theta}(\pi_{ijk}\mid s_{i},\mathcal{E}_{ij})),\\ &A^{(1)}_{ij}=R^{(1)}_{\lambda}(\mathcal{E}_{ij})-b_{\lambda}^{(1)}(s_{i}),\\ &A^{(2)}_{ijk}=R^{(2)}_{\lambda}(\pi_{ijk})-b^{(2)}_{\lambda}(s_{i},\mathcal{E}_{ij}).\end{split}$ (19)

### v4

# Beyond Simple Graphs: Neural Multi-Objective Routing on Multigraphs

@@ Abstract
Learning-based methods for routing have gained significant attention in recent years, both in single-objective and multi-objective contexts. Yet, existing methods are unsuitable for routing on multigraphs, which feature multiple edges with distinct attributes between node pairs, despite their strong relevance in real-world scenarios. In this paper, we propose two graph neural network-based methods to address multi-objective routing on multigraphs. Our first approach operates directly on the multigraph by autoregressively selecting edges until a tour is completed. The second model, which is more scalable, first simplifies the multigraph via a learned pruning strategy and then performs autoregressive routing on the resulting simple graph. We evaluate both models empirically, across a wide range of problems and graph distributions, and demonstrate their competitive performance compared to strong heuristics and neural baselines [^1] Our code will be released publicly upon paper acceptance..
@@ 1 Introduction
The field of neural combinatorial optimization has grown significantly in recent years and vehicle routing problems in particular have attracted much attention (Zhou et al., 2024a). While early works focused on the Traveling Salesman Problem (TSP) (Vinyals et al., 2015; Bello et al., 2017), new learning-based methods for vehicle routing can solve a wide range of problems efficiently and effectively, often surpassing classical state-of-the-art heuristics (Zhou et al., 2024b; Drakulic et al., 2025). Yet, existing methods have one limitation in common: they assume problems are defined on simple graphs. However, multigraph formulations, featuring several edges between each node pair, become relevant as soon as there are competing edges that cannot be chosen between a priori. Such situations typically occur when edges have more than one feature of interest, such as both travel time and distance.
In spite of the high practical relevance of multigraph formulations (Lai et al., 2016; Ben Ticha et al., 2017), current learning-based methods are incapable of handling them due to two main reasons. Firstly, many state-of-the-art neural solvers rely on transformers to encode the problem instance. While these work well in the Euclidean setting (Kool et al., 2018) and with some modifications on asymmetric, directed graphs (Kwon et al., 2021), they lack the capability to encode multigraph structures. Secondly and more importantly, planning routes in multigraphs requires both selecting the node order and which edges to traverse, making current decoding strategies unsuitable.
In this work, we aim to bridge the gap between learning-based methods for routing and accurate network representations given by multigraphs. We focus on the Multi-Objective (MO) setting, as several competing objectives naturally translates to several competing edges between each node pair. Nevertheless, our methods are general and can easily be extended to single-objective settings. Concretely, our main contributions are:
  • We design models to handle multigraph inputs. To our knowledge, we present the first neural solvers designed for such structures. In Appendix A, we argue further why current methods are insufficient, even when combined with graph transformations.
  • Our models are also capable of handling a variety of MO routing problems on asymmetric graphs. In contrast, current neural MO methods are either problem-specific or designed exclusively for the Euclidean setting, where edge costs are symmetric and determined only by the node coordinates. The non-Euclidean, asymmetric setting is generally viewed as more practically relevant (Boyacı et al., 2021).
@@
  • Experimentally, we show that both our approaches achieve competitive results for several routing problems on asymmetric graphs and multigraphs, including variants of the multi-objective TSP and multi-objective capacitated vehicle routing problem.
@@ 2.1 Learning for Routing
Motivated by the success of end-to-end learning in fields such as image classification and language modeling, neural solutions for routing have gained increasing attention during the last years. Broadly, methods can be classified by how they obtain tours. Two prominent alternatives are autoregressive construction, based on iterative building of tours one node at a time, and non-autoregressive construction, that outputs heatmaps in a one-shot fashion for down-stream decoding with some search method. Examples from the first category include Vinyals et al. (2015); Kool et al. (2018); Kwon et al. (2020). Examples from the second category are Joshi et al. (2019); Fu et al. (2021); Sun & Yang (2023). For a more comprehensive discussion of these methods, we refer the reader to Appendix B, which presents additional related work.
The literature on MO routing is more sparse than in the single-objective case. A key contribution was made by Lin et al. (2022), who proposed an autoregressive construction approach using a single neural model to learn the Pareto set of solutions. Regarding asymmetric problems, the only existing learning-based methods we are aware of are those introduced by Santiyuda et al. (2024) and Zhou et al. (2025). However, both these approaches are tailored to specific problems and have been evaluated within narrowly defined domains. In contrast, our method is capable of addressing a broad class of routing problems and is evaluated across various problems and settings.
@@ 2.2 Routing on Multigraphs
The multigraph representation has gained increasing attention in the operations research literature on Vehicle Routing Problems (VRPs) in recent years. Garaix et al. (2010) performed the first study in this regard and later works include those by Lai et al. (2016); Ben Ticha et al. (2017; 2019) and Tikani et al. (2021). A consistent finding is that the multigraph representation leads to considerably improved solutions compared to planning on simple graphs. For example, in the VRP with time windows, Ben Ticha et al. (2017) reported cost reductions of up to $10.5\%$ on real-world instances, while Lai et al. (2016) found average savings of around $5\%$ for a heterogeneous VRP. However, these benefits typically come at the cost of increased problem complexity. Learning-based methods can thus fill a clear need for new solvers in this area, as they have proven to be efficient, effective and broadly applicable. To the best of our knowledge, this work presents the first learning-based approach to address combinatorial routing problems such as the TSP and VRPs on multigraphs. In contrast, current studies either use traditional heuristics or exact solvers.
@@ 3 Problem Formulation
where $f:\Pi\to\mathbb{R}^{m}$ with $m>1$. Here, $\pi\in\Pi$ represents a feasible route in the multigraph $G$, encoded as a sequence of edges, and $f(\pi)$ is the associated cost vector.
@@
To obtain this set, we utilize a typical method: decomposition through scalarization. The idea is to translate the MO problem into several subproblems that can be solved separately. Let $\lambda\in\mathbb{R}^{m}$ such that $\lambda_{i}\geq 0\>\forall i$ and $\sum_{i}\lambda_{i}=1$ be a vector that controls the preference among the objectives. Then, under Linear Scalarization (LS), the single-objective cost is $f_{\lambda}(\pi)=\sum_{i}\lambda_{i}f_{i}(\pi)$ and the corresponding subproblem is $\min_{\pi}f_{\lambda}(\pi)$. Unfortunately, it is well-known that there may be solutions in the PS that do not solve an LS subproblem for any preference $\lambda$ (Ehrgott, 2005). The Chebyshev scalarization has more favorable properties. Let $z^{*}_{i}<\min_{\pi}f_{i}(\pi)$ be an ideal value for objective $i$. Then, the Chebyshev scalarized cost is
@@
Given a solution in the PS, there always exists a preference $\lambda$ and corresponding Chebyshev subproblem for which the solution is optimal (Choo & Atkins, 1983).
@@ 4 GNN-Based Multigraph Solver
Here, $||$ is the concatenation operation, $e_{l}$ denotes an edge embedding for edge $l$, $W_{1}^{\prime}$ and $W_{1}^{\prime\prime}$ are trainable weight matrices and $E^{+}(i)$, $E^{-}(i)$ denote outgoing and incoming edges respectively. New edge embeddings are then computed according to
@@
where $\text{start}(l)$ and $\text{end}(l)$ denote the start and end nodes of the directed edge $l$ and $W_{2}$ is another trainable matrix. The attention layers are arranged into a single GREAT-layer together with feed-forward layers, residual connections and normalization according to the original transformer architecture (Vaswani et al., 2017). Note that the residual connections are important as they allow parallel edges to have differing embeddings.
@@ 4.1 Edge-Based GMS
$p_{\theta(\lambda)}(\pi\mid s)=p_{\theta(\lambda)}(\pi_{1}\mid s)\prod_{t=2}^{T}\;p_{\theta(\lambda)}(\pi_{t}\mid\pi_{1:t-1},s).$ (9)
Here, $\theta(\lambda)$ is the weights of the model given the current preference $\lambda$. As Navon et al. (2021) and Lin et al. (2022), we utilize a Multi-Layer Perceptron (MLP) hyper-network to preference-condition the decoder, while keeping the encoder preference-agnostic. Consequently we form the decoder weights according to $\theta_{\text{dec}}(\lambda)=\text{MLP}(\lambda)$ and set $\theta(\lambda)=[\theta_{\text{enc}},\theta_{\text{dec}}(\lambda)]$. Regarding the decoder architecture, it is an edge-based variant of the Multi-Pointer (MP) decoder from Jin et al. (2023). We detail this component further in Appendix C.
[…]
Figure 1: Edge-based GMS and its most important components.
Note that a multigraph with $N$ nodes and $M$ edges between each pair of nodes has $\mathcal{O}(MN^{2})$ edges in total. Moreover, the decoder must be run $\mathcal{O}(N)$ times to construct a full tour and we roll out $\mathcal{O}(N)$ sample trajectories according to the POMO framework (Kwon et al., 2020). Hence, an edge-based decoder scales as $\mathcal{O}(MN^{4})$ compared to a node-based one that scales as $\mathcal{O}(N^{3})$. Consequently, we propose a second architecture that employs a node-based decoder and thereby scales better to higher node counts in terms of memory requirements and runtime.
@@ 4.2 Dual Head GMS
Again, the weights $\tilde{\theta}(\lambda)$ are generated by an MLP hyper-network. We detail the architecture for the selection-decoder further in Appendix C.
The edge choice from the first decoder is fed back into the encoder, where the embeddings of corresponding edges are aggregated into node embeddings using equation 7. After this final GREAT layer, we find it beneficial to add a small number of standard transformer layers to obtain more expressive node embeddings for the second decoder, which constructs routes.
The routing-decoder is a preference-conditioned version of the multi-pointer decoder (Jin et al., 2023). Similarly to the edge-based version above, given the partial tour $\pi_{1:t-1}$, this decoder outputs the probabilities $p_{\theta(\lambda)}(\pi_{t}\mid\pi_{1:t-1},s,\mathcal{E})$ for the next construction step, where $\pi_{t}$ now denotes a node and the probability depends on the edge choice $\mathcal{E}$.
@@
Figure 2: Dual head GMS and its most important components
@@ 4.3 Training Algorithm
Our first proposed model, GMS-EB, is trained using the Multi-Objective REINFORCE from Lin et al. (2022). Given a preference $\lambda\sim\Lambda$ and instance $s\sim S$, the model parameterizes a policy $p_{\theta(\lambda)}(\cdot\mid s)$ according to equation 9. A tour $\pi$ is associated with a Chebyshev reward corresponding to the negative scalarized cost equation 5:
@@
To stabilize the training, we form a baseline $b_{\lambda}(s_{i})$ by rolling out $K$ tours for the same instance $s_{i}$ and averaging the reward, according to the POMO framework (Kwon et al., 2020). The policy gradient is then utilized to update the model weights and it is approximated, given one preference $\lambda$, as
$\nabla J(\theta)\approx\frac{1}{BK}\sum_{i,j=1}^{B,K}\left(R_{\lambda}(\pi_{ij})-b_{\lambda}(s_{i})\right)\nabla_{\theta}\log(p_{\theta(\lambda)}(\pi_{ij}\mid s_{i})),$ (13)
where $B$ is the batch size of instances.
For the second model, the training is more involved. Given an edge selection $\mathcal{E}\sim q_{\tilde{\theta}(\lambda)}$ from the selection-head according to equation 10, the routing-head parameterizes a policy $p_{\theta(\lambda)}(\cdot\mid s,\mathcal{E})$ and obtains a scalar reward $R^{(2)}_{\lambda}(\pi)=R_{\lambda}(\pi)$ by equation 11 as usual. With a fixed $\tilde{\theta}$, the objective for this head is to maximize
@@
[…]
@@
$\begin{split}&\nabla J_{1}(\tilde{\theta})\approx\frac{1}{BK_{1}}\sum_{i,j=1}^{B,K_{1}}\left(R^{(1)}_{\lambda}(\mathcal{E}_{ij})-b_{\lambda}^{(1)}(s_{i})\right)\nabla_{\tilde{\theta}}\log(q_{\tilde{\theta}}(\mathcal{E}_{ij}\mid s_{i})),\\ &\nabla J_{2}(\theta)\approx\frac{1}{BK_{1}K_{2}}\sum_{i,j,k=1}^{B,K_{1},K_{2}}\left(R^{(2)}_{\lambda}(\pi_{ijk})-b^{(2)}_{\lambda}(s_{i},\mathcal{E}_{ij})\right)\nabla_{\theta}\log(p_{\theta}(\pi_{ijk}\mid s_{i},\mathcal{E}_{ij})).\end{split}$ (17)
@@
In addition, we utilize a form of curriculum learning to speed up training for both models. We start by training on small graph sizes and then gradually increase the problem size. For GMS-DH in the multigraph setting, we also start by training the routing-head on simple graphs to ensure that the approximation equation 16 is accurate even during the initial epochs of multigraph training. Further details regarding the curriculum learning can be found in Appendix D.
5 Experimental Studies